### v1

# No-exclaves percolation on random networks

Abstract
No-exclaves percolation (NExP) is a nonlocal percolation process in which the components are formed not only by the connected occupied nodes but also by the agglomeration of empty nodes completely surrounded by the occupied nodes. It has been studied in low dimensions, displaying such novel phenomena as the discontinuous transition to complete percolation. However, its characteristics in complex networks are still unexplored. In this paper, we study the NExP on random networks by developing mean-field solutions using the generating function formalism. Our theory allows us to determine the size of the giant no-exclaves component as well as the percolation threshold, which are in excellent agreement with Monte Carlo simulations on random networks and some real-world networks. We show that on random networks NExP exhibits three phases and two transitions between them: the phases are characterized by the presence or absence of not only the giant no-exclaves component but also the giant unoccupied component, which is the giant connected component composed solely of unoccupied nodes. This work offers theoretical understanding on the anatomy of phase transitions in the NExP process.
keywords:
No-exclaves, Percolation, Network Robustness
[…]
1 Introduction
Understanding percolation of complex networks has an impact on our ability to predict and control the behaviors of complex systems as it can provide essential information regarding the connectivity and robustness of complex systems stauffer ; error_attack ; newman_book ; dlee . Classical percolation models are based on the notion of connectivity in which a pair of nodes are considered a part of the same connected component if there exists at least one connected path between them stauffer . This model has provided a simple yet powerful theoretical framework across various phenomena, including epidemic spreading newman02 ; pastor02 , traffic flows transport , the stability of power grid power ; yang ; chilean , and the resilience of networks error_attack ; callaway ; cohen2000 . There have also been numerous studies on variants of the ordinary percolation models on networks that have modified the concept of connectivity such as $k$-core percolation kcore , bootstrap percolation bootstrap , $k$-selective percolation kim_goh , and mutual percolation buldyrev ; son ; baxter12 ; min14 .
One of the most important applications of percolation theory in complex networks is the robustness of networks against random failures or attacks error_attack ; callaway ; cohen2000 ; cohen2001 ; holme . The robustness of networked systems is often measured by how the size of the largest connected component responds when some nodes in networks are failed error_attack ; holme . The premise underlying the measurement is that it is essential for nodes to maintain their connectivity in order to function properly error_attack ; buldyrev . Additionally, when nodes become non-functioning due to failure or detachment from the largest component these nodes are assumed to be unable to recover permanently in most studies on the robustness of networks error_attack ; newman_book . In reality, however, damaged nodes can often recover with communications to properly functioning nodes that surround the damaged nodes min14 ; recovery ; shekhman16 ; shang ; wu2022 . For instance, temporary breakdown in real-world networks such the brain and financial systems can recover as long as the damage is localized.
The no-exclaves percolation (NExP) is a model of incorporating such a recovery rule that components of failed nodes that are completely surrounded by active (unfailed) nodes recover to be unfailed and merge into the surrounding unfailed component nexp . The failed components that are surrounded by unfailed nodes are referred to as “exclaves” which is the terminology in political geography and means a part of a district geographically isolated from the mainland by surrounding alien territories. Our model does not allow for exclaves, hence the name “no-exclaves percolation,” or NExP for short. After applying the no-exclaves rule for the recovery, we identify connected components and assess the robustness of networks using the concept of the giant NExP component. The no-exclaves rule in our model is closely related to the no-enclaves percolation on two dimensions offering a theoretical explanation for a motor-driven collapse of cytoskeletal systems sheinman ; gunnar ; reply ; hu2016 . Note that the no-enclaves rule does not readily apply to complex networks; the no-exclaves rule, however, is applicable to generic networks, while keeping the ingredient of nonlocality intact. Furthermore, the no-exclaves rule implies long distance communications between active nodes in some sense castellano2020 . From this perspective, it might have potential applicability to quantum communication and quantum network issues as well meng ; extended .
It is known from numerical simulations that in low-dimensional systems the no-exclaves rule significantly affects the behavior of percolation transitions in spatial networks and Euclidean lattices nexp . However, there is still a lack of the studies for NExP on networks, as well as the analytical understanding of NExP process. The main aim of the study is to make progress in filling this gap by developing the exact mean-field solutions to NExP based on generating function methods in order to address the resilience of complex networks. We derive mean-field solutions to NExP on random networks, leading to an analytical insight to the effect of the no-exclaves rule and the anatomy of novel phase transitions displayed by the NExP process.
The paper is organized as follows. In Sec. II, we introduce the no-exclaves percolation model and explain how to apply no-exclave rules. We also introduce the concept of the giant NExP component as the order parameter. Next, we derive an analytical theory by using generating function formalism, to compute the size of giant NExP component and the average NExP component size in random networks (Sec. III). Based on the theory, we identify and analyze the location of percolation transitions. Finally, we apply our theory to random regular networks, Erdős-Rényi networks, and real-world networks from empirical data (Sec. IV). The summary and discussions are presented in Sec. V.
2 No-Exclaves Percolation Model
In this section, we present the no-exclaves percolation on networks, which was originally studied for lattices nexp . The NExP process is defined as follows. Initially, all nodes are in an empty state in a given network. With the occupation probability $q$, we set each node to be occupied, randomly and independently. In terms of the network robustness, $1-q$ corresponds to failure probability. So far, the procedures are the same as the usual site percolation problem. Next, we identify the exclaves, which are the components composed of unoccupied nodes that are completely surrounded by the occupied nodes, except for the largest such one. Finally, the exclave gets merged into a bigger component comprising itself and all the neighboring occupied components. Note that during this process multiple connected components, and even multiple exclaves, can merge into a single, larger component. We refer to the resulting combined components as the NExP components. We then measure the size $\Phi$ of the largest NExP component as an order parameter and use occupation probability $q$ as a control parameter. The size $\Phi$ is equivalent to the probability that a randomly selected node belongs to the largest NExP component.
Figure 1 shows an example of no-exclaves percolation. Each node in a network is occupied with occupation probability $q$ [marked blue in Fig. 1(a)] or empty with probability $1-q$ [marked white in Fig. 1(a)]. Then, we identify the exclaves red-shaded in Fig. 1(b), the components of connected empty nodes that are completely surrounded by occupied nodes. Practically, if all links leading out from the components composed of empty nodes connect to occupied nodes, the components correspond to exclaves. Note that the “giant” connected component composed of empty nodes is not an exclave. Next, the exclave components merge with the neighboring occupied components, becoming “virtually”-occupied themselves as shown in Fig. 1(c). After agglomerating all the exclaves and surrounding occupied components, we identify the NExP component, the connected component of both real- and virtually-occupied nodes [Fig. 1(d)]. The largest NExP component in (d) is in general greater than the largest connected component in (a).
Figure 1: Schematic illustrations of NExP on a network. (a) Each node in the network is either occupied (blue) with a probability $q$, or empty with a probability $1-q$. (b) Exclaves (red), components of empty nodes surrounded by occupied nodes, are identified. (c) The exclaves merge with surrounding components. (d) The giant NExP component is identified, marked by the grey shaded area.
3 Generating Function Methods
In this section, we derive the analytic solution of NExP on random graphs using generating function formalism, applicable to locally-treelike networks in the thermodynamic limit. We observe that according to the no-exclaves rule, there can be three categories of components in NExP on large networks: i) a giant NExP component, ii) small (non-giant) NExP components, and iii) a giant unoccupied component. A giant unoccupied component, which contains a finite fraction of nodes in the thermodynamic limit and extends to infinity, cannot be completely surrounded and thus is not an exclave. Note that we include isolated unoccupied components within the NExP component as there is no need to distinguish between them for our purposes.
Based in this observation, our strategy for the analytical solutions of NExP on networks is the following: We first identify the giant unoccupied component that should be excluded when obtaining NExP components by using a standard technique for the site percolation newman . The major difference is that we here consider a giant “unoccupied” component rather than occupied one that has been normally focused on. Then we derive self-consistency equations for the probability that a node arrived by following a randomly chosen link belongs to small NExP component. The size of the giant NExP component, denoted as $\Phi$, is determined by the complementary probability to the sum of the giant unoccupied component and the small NExP components.
3.1 The size of giant unoccupied component
As noted above, we first calculate the size $S_{0}$ of a giant unoccupied component in order to obtain the size $\Phi$ of the giant NExP component. The size $S_{0}$ can be obtained by following a conventional technique for the site percolation problem newman . It is important to note that since we calculated the size of the giant unoccupied component, we determine the giant component among “unoccupied” nodes, rather than among occupied nodes as in ordinary site percolation problems.
Consider a random network with its degree distribution $P(k)$. We define the degree and excess degree generating functions
${G}_{0}(x)$ $=\sum_{k=0}^{\infty}P(k)x^{k},$ (1)
${G}_{1}(x)$ $=\sum_{k=1}^{\infty}\frac{kP(k)}{z}x^{k-1},$ (2)
where the excess degree stands for the degree for a node reached by following a randomly chosen link newman , and $z$ stands for the mean degree, $z=\langle k\rangle=\sum_{k}kP(k)$. Let $R(s)$ be the probability that a randomly chosen node belongs to a small (non-giant) unoccupied component with size $s$. Here an unoccupied component stands for a connected component composed of solely unoccupied nodes. Then we define a generating function of $R(s)$ as ${H}_{0}(x)=\sum_{s=0}^{\infty}R(s)x^{s}$. We also define a generating function ${H}_{1}(x)$ that generates the size of a small unoccupied component reached by following a randomly chosen link.
On locally tree-like networks, the size of the small unoccupied component to which node $i$ reached by following a randomly chosen link belongs is either zero if node $i$ is occupied (with probability $q$) or the sum of the sizes of the unoccupied components to which node $i$’s neighbors belong plus one for node $i$ itself if node $i$ is unoccupied (with probability $1-q$). Therefore, ${H}_{1}(x)$ and $H_{0}(x)$ satisfy the following self-consistency equations
${H}_{1}(x)$ $=q+(1-q)x{G}_{1}({H}_{1}(x)),$ (3)
${H}_{0}(x)$ $=q+(1-q)x{G}_{0}({H}_{1}(x)).$ (4)
We define $w$ as the probability for a node arrived at following a randomly chosen link does not belong to a giant unoccupied component. According to the definition of $H_{1}(x)$, the value of $w$ is identical to $H_{1}(1)$, as given by Eq. (3)
$w=q+(1-q){G}_{1}(w).$ (5)
In addition, as given the definition ${H}_{0}(x)$, we have the total probability that a randomly chosen node belongs to a small unoccupied component as ${H}_{0}(1)=\sum_{s=0}^{\infty}R(s)$. Note that it is not necessarily normalized such that ${H}_{0}(1)=1$ because there may exist a giant unoccupied component that is not included in $R(s)$. The size of the giant unoccupied component, $S_{0}$, on a random network is equal to the probability that a randomly chosen node belongs to the giant unoccupied component. In conclusion, once we obtain $w$, the size $S_{0}$ of the giant unoccupied component can be obtained as
$S_{0}=(1-q)\left[1-G_{0}(w)\right].$ (6)
3.2 Self-consistency equations for NExP component sizes
In this section, we attempt to derive the self-consistency equations for NExP component sizes. Let $Q(\phi)$ be the probability that a randomly chosen node belongs to a finite (non-giant) NExP component with size $\phi$ excluding both the giant NExP component and the giant unoccupied component. Then we define the generating function of $Q(\phi)$, represented as $F_{0}(x)=\sum_{\phi=0}^{\infty}Q(\phi)x^{\phi}$. We also define the generating function ${F}_{1}(x)$ that generates the size of finite NExP components for a node reached by a randomly chosen link. Note that $Q(\phi)$ may not be normalized to be unity since it excludes both a giant NExP component and a giant unoccupied component.
We then attempt to derive self-consistency equations of ${F}_{0}(x)$ and ${F}_{1}(x)$ on a locally tree-like network. Assume that if we arrive at node $i$ which belongs to a small NExP component by following a randomly chosen link, the size $\phi$ of the small component including node $i$ is one for the node $i$ plus the sum of the sizes of the small components to which node $i$’s neighbors belong. A key point to consider is that the probability of belonging to a size $\phi$ component can vary depending on whether node $i$ is i) occupied or ii) unoccupied.
i) If node $i$ is occupied, a neighbor of node $i$ can be a part of either a small NExP component or a giant unoccupied component. Note that if a neighbor belongs to a giant unoccupied component, it contributes by zero to the small component’s size. Then, the probability that a neighbor of an “occupied” node belongs to a small component with size $\phi$ is generated by $(1-w)+{F}_{1}(x)$. There is an additional $(1-w)$ term with zeroth order in $x$ because it represents a probability for a giant unoccupied component, which is counted as if a zero-size NExP component.
ii) On the other hand, if node $i$ is unoccupied, every neighbor of node $i$ cannot be a part of a giant unoccupied component. If this were the case, node $i$ would naturally belong to a part of the giant unoccupied component. Therefore, the probability that a neighbor of an “unoccupied” node belongs to a small component with size $\phi$ is generated by ${F}_{1}(x)$
Summarizing these possibilities, we arrive at the following self-consistency equations for ${F}_{0}(x)$ and ${F}_{1}(x)$:
$F_{1}(x)$ $=qxG_{1}(1-w+F_{1}(x))+(1-q)xG_{1}(F_{1}(x)),~{}$ (7)
$F_{0}(x)$ $=qxG_{0}(1-w+F_{1}(x))+(1-q)xG_{0}(F_{1}(x)).~{}$ (8)
We define $u$ as the probability that a node by following a randomly chosen link belongs to a small NExP component. This means it belongs to neither the giant NExP component nor the giant unoccupied component. The probability $u$ is identical to $F_{1}(1)$ according to the definition of $F_{1}(x)$. The value of $u$ can be obtained through numerical iteration of Eq. (7), as given by
$u=q{G}_{1}(1-w+u)+(1-q){G}_{1}(u),$ (9)
with $w$ from Eq. (5).
3.3 The size of giant NExP component
The size $\Phi$ of a giant NExP component on a random network is identical to the probability that a randomly chosen node belongs to the giant NExP component. The complementary probability of $\Phi$ is that a randomly chosen node either belongs to a small NExP component, with probability $F_{0}(1)$, or belongs to a giant unoccupied component, with probability $S_{0}$. The probability $F_{0}(1)$ can be easily expressed as
${F}_{0}(1)=q{G}_{0}(1-w+u)+(1-q){G}_{0}(u),$ (10)
according to Eq. (8). Then, we finally arrive at the following expression for the size $\Phi$ of the giant NExP component as
$\Phi$ $=1-{F}_{0}(1)-S_{0}$ (11)
$=1-q{G}_{0}(1-w+u)-(1-q)[1-{G}_{0}(w)+G_{0}(u)],$
with $w$ from Eq. (5) and $u$ from Eq. (9).
In summary, our strategy for obtaining the size $\Phi$ of the giant NExP component has been as follows. We first determine the probability $w$ that a node arrived by following a link does belong to the giant unoccupied component by solving Eq. (5), and subsequently the size $S_{0}$ of the giant unoccupied component by Eq. (6). In addition, we obtain the probability $u$ that a node arrived by following a link belongs to a small NExP component by Eq. (9). Putting all together, we can obtain the size $\Phi$ of the giant NExP component by Eq. (3.3).
3.4 The average NExP size and percolation threshold
We derive expressions for the average size of small NExP components, excluding the giant NExP component. When the giant NExP component does not exist, a node reached by following a randomly chosen link must belong to either a small component with probability $u$ or a giant unoccupied component with probability $1-w$, resulting in the condition $11-w+u=$ 1. It simplifies to the condition $u=w$. With this condition, the average small NExP component size can be expressed as
$\chi$ $=\left.\frac{d{F}_{0}(x)}{dx}\right|_{x=1}$
$=F_{0}(1)+[qG_{0}^{\prime}(1)+(1-q)G_{0}^{\prime}(w)]F_{1}^{\prime}(1),$ (12)
where ${}^{\prime}$ denotes the derivative with respect to its argument $x$. From Eq. (7), we obtain
$F_{1}^{\prime}(1)=\frac{w}{1-qG_{1}^{\prime}(1)-(1-q)G_{1}^{\prime}(w)}.$ (13)
Since we consider the case when the giant NExP component does not exist, the average component size is given by
$\chi=F_{0}(1)+\frac{w[qG_{0}^{\prime}(1)+(1-q)G_{0}^{\prime}(w)]}{[1-qG_{1}^{\prime}(1)-(1-q)G_{1}^{\prime}(w)]}.$ (14)
The percolation threshold $q_{c}$ where a giant NExP component first appears is located at the point where the average component size diverges, satisfying
$q_{c}G_{1}^{\prime}(1)+(1-q_{c})G_{1}^{\prime}(w)=1.$ (15)
The percolation threshold $q_{c}$ of the giant NExP component is in general not amenable to compact expression in terms of $z$ or other moments of $P(k)$ because $w$ also contains $q$, yet can be computed numerically. Beyond this threshold $q_{c}$, the size of the giant NExP component continues to grow as more occupied nodes are added.
In NExP, there is another transition where a giant unoccupied component disappears as $q$ increases. In a scenario where a giant unoccupied component vanishes, there is a potential for all nodes within a connected component to be part of the giant NExP component. We refer the point of this transition as the “complete” percolation point, denoted as $q^{*}$. The complete percolation point $q^{*}$ corresponds to the location in which the size $\Phi$ of the giant NExP component becomes unity if all nodes of the underlying network belong to a single connected component. The location of $q^{*}$ can be identified by the divergence of the average size $\langle s\rangle$ of unoccupied components. The condition for $q^{*}$ is given from Eq. (3) as
$G_{1}^{\prime}(1)=\frac{1}{1-q^{*}}.$ (16)
Since $G_{1}^{\prime}(1)=\frac{\langle k^{2}\rangle-\langle k\rangle}{\langle k\rangle}$ for random networks, $q^{*}$ can be expressed compactly as
$q^{*}$ $=1-\frac{\langle k\rangle}{\langle k^{2}\rangle-\langle k\rangle}=\frac{\langle k^{2}\rangle-2\langle k\rangle}{\langle k^{2}\rangle-\langle k\rangle},$ (17)
which is directly related to the well-known result for the percolation on random networks callaway ; cohen2000 ; molloy1998 .
The transitions can be interpreted by the stability analysis of the solutions of Eqs. (5) and (9). The complete percolation point $q^{*}$ corresponds to the point where the trivial fixed point of $w$ in Eq. (5) becomes unstable. The trivial solution of Eq. (5) is $w=1$ since $G_{1}(1)$ is normalized to be $1$. The solution $w=1$ corresponds to no giant unoccupied component considering the meaning of $w$. That means the giant unoccupied component disappeared at $q^{*}$. On the other hand, the NExP threshold $q_{c}$ can be obtained by the stability analysis of Eq. (9). The trivial solution of Eq. (9) is $u=w$ because Eq. (9) reduces into Eq. (5) when $u=w$, or equivalently $11-w+u=$ 1. The solution $u=w$ represents no giant NExP component and it becomes unstable at $q_{c}$. Therefore another solution becomes stable, leading to the emergence of the giant NExP component.
These two transition points $q_{c}$ and $q^{*}$ do not in general coincide on random networks. The percolation transition $q_{c}$ is generally smaller than the complete point $q^{*}$. Therefore, there can be three phases in NExP, depending on the existence and extinction of the giant NExP component and the giant unoccupied component: We have phases in which i) the giant unoccupied component exists but giant NExP component does not when $q<q_{c}$; ii) both giant NExP component and giant unoccupied component exist when $q_{c}<q<q^{*}$; iii) the giant NExP component exists but giant unoccupied component does not when $q>q^{*}$. The multiple percolation transitions in network NExP are in stark contrast to the 2D NExP in which the two occur at the same point nexp . Our solutions suggest that the discontinuous percolation transition into complete percolation displayed by 2D NExP can be resolved into two transitions combined into one.
4 Results
4.1 On Random Regular Networks
To gain insight on the solutions of NExP, we derive explicit expressions of our theory applied to random regular (RR) networks with degree $z$ and a uniform occupation probability $q$. The degree distribution of RR networks is a Kronecker delta distribution, $P(k)=\delta_{k,z}$. In RR networks, the probability that a node reached by following a link belong to a finite unoccupied component $w$ and belong to a finite NExP component $u$ are given by the coupled self-consistency equations as
$w$ $=q+(1-q)w^{z-1}$ (18)
$u$ $=q(1-w+u)^{z-1}+(1-q)u^{z-1}.$ (19)
For an explicit example of $z=3$, the solutions of $w$ are $1$ and $q/(1-q)$, which correspond to the phase without a giant unoccupied component and with a giant unoccupied component, respectively. The trivial fixed point $w=1$ corresponding to no giant unoccupied component becomes unstable when $q<q^{*}$ where $q^{*}$ is the complete percolation point. The value of complete percolation point between the two phases is given by
$q^{*}=\frac{z-2}{z-1}=\frac{1}{2}.$ (20)
Depending on $q$, the value of $w$ is explicitly given by
$w=\left\{\begin{array}[]{ll}\frac{q}{1-q},&q<q^{*},\\ 1,&q>q^{*}.\end{array}\right.$ (23)
While there is the giant unoccupied component when $q<q^{*}$, the giant unoccupied component disappears when $q>q^{*}$.
When $w=1$ for $q>q^{*}$, $u$ becomes 00. If $q<q^{*}$ indicating that $w=q/(1-q)$, the possible solutions of $u$ are either $u=q/(1-q)$ and $(1-2q)^{2}/(1-q)$. The solution $u=w=q/(1-q)$ corresponds to the case when there is no giant NExP component. If $u=(1-2q)^{2}/(1-q)$, $1-w+u$ becomes less than unity meaning that there is a giant NExP component. The percolation threshold between two phases is given by
$w(q_{c})$ $=\frac{1}{z}=\frac{1}{3}.$ (24)
$q_{c}$ $=\frac{1}{4}.$ (25)
Depending on $q$, the value of $u$ is given by
$u=\left\{\begin{array}[]{ll}\frac{q}{1-q},&q<q_{c},\\ \frac{(1-2q)^{2}}{1-q},&q_{c}<q<q^{*},\\ 0,&q>q^{*}.\end{array}\right.$ (29)
Therefore we have three regimes of no-exclaves percolation in RR networks: i) a giant NExP component does not exist but a giant unoccupied component exists ($q<q_{c}$); ii) both giant NExP component and giant unoccupied component exist ($q_{c}<q<q^{*}$); and iii) an only giant NExP component whose size is unity exists ($q>q^{*}$). Since we can express $u$ and $w$ in closed forms for RR networks with $z=3$, we can explicitly obtain the expression for the size of the giant NExP component $\Phi$ as:
$\Phi=\left\{\begin{array}[]{ll}0,&q<q_{c}~{},\\ \frac{(4q-1)(8q^{3}-6q^{2}+1)}{1-q},&q_{c}<q<q^{*}~{},\\ 1,&q>q^{*}~{}.\end{array}\right.$ (33)
As shown in Fig. 2, the probabilities $w$ and $u$, and size of the giant NExP component $\Phi$ as a function of the occupation probability $q$ show perfect agreement with the Monte Carlo simulation results.
Figure 2: Analytic predictions of $\Phi$, $w$, and $1-w+u$ with respect to $q$ on RR networks with $z=3$ and $N=10^{6}$ are shown. Theoretical results (lines) and numerical results (points) are shown together. Error-bars denote the standard deviation from $10^{6}$ independent runs.
4.2 On Erdős-Rényi networks
We consider the no-exclaves percolation model on Erdős-Rényi (ER) networks, which have a Poisson degree distribution $P(k)=e^{-z}z^{k}/k!$ in the large $N$ limit, where $z$ is the average degree. The generating functions for ER networks are given by $G_{0}(x)=e^{z(x-1)}$ and $G_{1}(x)=e^{z(x-1)}$. This leads us to the coupled equations for the variables $w$ and $u$, given by
$w$ $=q+(1-q)e^{z(w-1)}~{},$ (34)
$u$ $=qe^{z(u-w)}+(1-q)e^{z(u-1)}.$ (35)
The complete percolation point and the percolation threshold for NExP can be determined by Eqs. (17) and (15), respectively. These conditions show the instability of the trivial solutions, $w=1$ for the complete percolation point and $u=w$ for the percolation threshold. For ER networks, the specific values for these thresholds are given by
$q^{*}=1-\frac{1}{z},\quad w(q_{c})=\frac{1}{z}.$ (36)
We have three regimes in the NExP model, similar to RR networks. i) When $q<q_{c}$, there is no giant NExP component but a giant unoccupied component exists. ii) When $q_{c}<q<q^{*}$, both a giant NExP component and a giant unoccupied component exist. iii) When $q>q^{*}$, only a giant NExP component exists. Furthermore, when $q>q_{c}$, the size of the giant NExP component $\Phi$ can be expressed as
$\Phi$ $=1-qe^{z(u-w)}-(1-q)[1-e^{z(w-1)}+e^{z(u-1)}].$ (37)
where $w$ and $u$ can be obtained by Eqs. (34-35).
We compare the theory derived using the generating function method with Monte Carlo simulations on ER networks with average degree $z=3/2,2,3$ and network size $N=10^{6}$. Figure 3 shows the size of the largest NExP component, $\Phi$, obtained from our theory (line) and numerical simulations (points) as a function of $q$. We observe excellent agreement between our theoretical predictions and the simulation results. Furthermore, we validate the locations of the percolation transition, $p_{c}$, and the complete transition point, $p^{*}$, which were predicted from Eq. (36).
Figure 3: The size of the giant NExP component $\Phi$ as a function of the occupation probability $q$ on ER networks with mean degree $z=3/2,2$, and $3$ and $N=10^{6}$. Numerical results (points) averaged over $10^{6}$ independent runs and analytic curves (lines) are shown together.
4.3 On real-world networks
We test our theory for the no-exclaves percolation on top of real-world networks built from empirical dataset. We use the two datasets: the Internet of autonomous system level caida and the U.S. airport network in 2010 obtained from “Open flights” flights . The router network of Internet represents that the traffic flow exchanges with its neighboring peers at autonomous system level, collected on November 5, 2007. The U.S. airport network represents the connection from the origin to destination airport. For the sake of simplicity, we ignore directionality and weights of the links from the data when we construct the networks. The router network contains $26,464$ nodes with $z=3.26$ and the airport network has $1,573$ nodes with $z=30.39$.
The theoretical predictions and numerical results of the size of the giant NExP component are shown in Fig. 4. We confirm that our theory generates a reliable prediction in the size of the giant NExP component for both the Internet [Fig. 4(a)] and the airport network [Fig. 4(b)]. We also measure the size of the giant connected component for random site percolation (denoted in short as RSP) for comparison callaway ; cohen2000 . We found that $\Phi$ for NExP is higher, compared to the size of giant component by standard node removal scenarios. It concludes that real-world networked systems can be more robust than expected when considering the recovery of exclaves. It is noteworthy that the shape of $\Phi$-curve for the real-world networks is different from those of Fig. 3: It does not display the two-transition character clearly. This is due mainly to the fact that both the real-world networks are “scale-free,” having broad distribution of degrees, driving $q_{c}$ towards zero and $q^{*}$ towards unity. Finite-size effect can also be attributed for the smearing of the transitions in real-world networks. Even so, our theory works remarkably well on these real-world networks.
Figure 4: The size $\Phi$ of the giant NExP component of two real-world networks: (a) the Internet’s router network of autonomous system level and (b) the U.S. airport network. Monte-Carlo simulation results (symbols) averaged over $10^{4}$ independent runs and analytic curves (lines) are shown together.
5 Summary and discussion
In this study, we have focused on the no-exclaves percolation (NExP) process on random networks and formulated a generating function approach to derive analytical solutions for this model. We identify the percolation threshold and complete transition point for the emergence of the giant NExP component and disappearance of the giant unoccupied component, respectively. In addition, we apply the theory to random regular graphs, Erdős-Rényi graphs, and real-world networks, and confirm the great agreement between our theoretical predictions and Monte Carlo simulations.
From theoretical perspective, our solution offers theoretical understanding for the multiple phase transition structure of NExP on random networks. It shows how the two-threshold behavior associated with the emergence of the giant NExP component and the extinction of the giant unoccupied component emerge as the occupation probability $q$ increases in network NExP. It also suggests that the single discontinuous percolation in 2D NExP nexp may be resolved into two transitions converging at the same point.
From the perspective of network robustness, our study implies that the recovery processes according to the no-exclaves rule can effectively mitigate the impact of random failures of nodes. When we remove nodes with failure probability $1-q$ as random failure problems error_attack ; cohen2000 the giant “occupied” component disappears at $q^{*}$. However, by introducing the no-exclaves rule, we can maintain the giant NExP component at the point $q_{c}$, which is in general less than $q^{*}$. In addition, we also found that the random failure above the complete percolation point $q^{*}$ does not make any effect in the no-exclaves percolation. In this sense, our study suggests an effective way to improve the network robustness with the recovery of failed exclaves.
Acknowledgments
This work was supported in part by the National Research Foundation of Korea (NRF) grants funded by the Korea government (MSIT) (No. 2020R1A2C2003669 (K-IG) and No. 2020R1I1A3068803 (BM)).